sammixboo (sammixboo):

Henry grabs one shoe at a time at random without replacement from a box of fourteen shoes in his closet. The box contains 6 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be brown? Round to the nearest percent if necessary.

sammixboo (sammixboo):

A. 6% B. 7% C. 16% D. 18%

OpenStudy (anonymous):

\[\frac{6}{14}\times \frac{5}{13}\] is my guess

sammixboo (sammixboo):

that is wat i thought to do

sammixboo (sammixboo):

so is tould be 15/91 which is rounded to 16% rigth